# custom.bib

### v1

@@ 4.3 Ablation study & Analysis
The motivation of including the zero-gist mechanism is to ensure that context-independent tokens can bypass the influence of the edit contexts. To assess the impact of zero-gist, we train a model without this mechanism and evaluate it on DUNE [dune] and MQuAKE [mquake] (see the “w/o zero-gist” line in Table 4). The results show a notable performance drop, particularly in multi-hop scenarios, suggesting that the cross-attention calculations may sometimes interfere with ordinary generation and our zero-gist strategy can mitigate this issue by allowing tokens to “attend to nothing”.
@@
We present the reason for our decision to use the KV cache from the second half of the model layers. To investigate this, we train a model using the KV cache from all layers and evaluate it on 1000 instances from ZsRE [zsre]. We record the probabilities allocated to the zero-gist in the cross-attention modules, as shown in Fig. 4b. The result shows that the zero-gist probabilities in layers 7-15 are generally lower than those in layers 1-6, and there is a notable drop in the zero-gist probability at layer 7. This suggests that, even when trained to use the full-model KV cache, the model mainly relies on information from deeper layers, since higher probabilities of the zero-gists indicate lower utilization of the actual edit contexts. A possible explanation is that more information is accumulated in the deeper layers, which aids both compression and selection processes. To verify our analysis, we also test the full layer trained model (see the “w/ full model” line in Table 4) under a batch editing setting, and observe a general decrease across all metrics. Additionally, restricting the KV cache to only the second half of the model could provide efficiency benefits with lower memory and computation costs.
@@
Applying a small temperature to the gist cross-attention sharpens the probability distribution over the gist KV caches, which facilitates the model’s ability to retrieve the correct information. We determine the appropriate temperature based on entropy, which has been shown to be an important factor in attention mechanisms [atten-entropy]. Specifically, we aim to keep the cross-attention entropy close to its optimal value, which occurs when it only needs to attend to one edit. To achieve this, we select 1000 instances from ZsRE [zsre] and calculate the entropy of the edit batch size 1. We then calculate the entropy for larger edit batch sizes (10, 20, 40, 60, 80, 100, 200, 300) and find the temperatures that align their entropy with the optimal case via gradient descent  [^7] Note that all entropy is calculated in a way that the query and answer part are just a copy of the context. We report the calculated temperature in Figure 4a. As expected, the temperature decreases as the edit batch size increases, but interestingly, it gradually converges to a specific value. Specifically, layers 9, 10, and 14 finally converge to around 0.5. To encourage more decisive selection, we slightly lower this value and set the temperature to T=0.45.
@@ 4.4 Efficiency Analysis
Finally, we present an efficiency analysis of different editing methods. Compared to many traditional editing methods that require model backward calculation, our method only requires one single forward pass for each editing context. In comparison to ICL, which needs to encode the entire concatenated edit context, our approach enables parallel encoding of multiple edits, leading to efficiency gains. Table 5 reports the time required to perform 100 edits for each method. Our method is obviously better than traditional editing methods and further improves over ICL by 17% to 27%. Additionally, compared to ICL, our approach only needs to maintain the KV cache of the gist representations from the deeper half layers, resulting in substantially lower memory cost.
Model InComeS ICL FT-M LoRA [LoRA] MEMIT [memit] EMMET [emmet] Llama-3.2-1B 0.69 0.95 6.42 46.54 200.28 192.65 Qwen2.5-7B 0.92 1.12 38.61 173.18 619.48 697.47
Table 5: Measured time (seconds) for 100 edits from COUNTERFACT [memit].
@@ 5 Related Work
The area of knowledge editing (or model editing) has experienced a thriving development in recent years. Researchers have explored various directions in this area. One typical direction is to adopt external memory for the edits. The memory formats applied by different researchers are diverse. Methods like SERAC [serac], IKE [ike], MeLLo [mello] adopt explicit non-parametric memory, which stores specific edit instances, and a retriever that is responsible for recalling relevant edits from the memory. For example, IKE uses KNN, and SERAC applies a trained classifier. Another line of work applies implicit parametric memory. CaliNET [calinet], T-Patcher [t-patcher] embeds the knowledge into a fixed number of neurons and adds them to the model. GRACE [grace] adopts a discrete key-value codebook with the value optimized for the desired knowledge. MELO [melo] applies dynamic LoRA blocks and indexes them via an internal vector database. KE [ke], MEND [mend] train a separate meta-model for editing. Another popular direction is to merge knowledge into the model directly. Methods like KN [kn], ROME [rome], R-ROME [r-rome], MEMIT [memit], PMET [pmet], CoachHooK [coachhook], and AlphaEdit [alphaedit] perform editing by tweaking the located FFN part of the model directly. However, some studies reveal that these methods could potentially bring about side effects in the original model [hurt-general-ability, Ocean-with-a-Spoon], leaving the real effectiveness of these methods to be further investigated.
@@ 6 Conclusion
In this paper, we propose InComeS, a scalable model editing method that integrates compression and selection mechanisms directly into the LLMs. InComeS adopts a context compression technique to condense the editing context to KV representations on top of the introduced gist tokens and takes advantage of the compressed KVs to efficiently retrieve the relevant editing context information. Experiments on four different and complex editing settings demonstrate the superiority of our method for comprehensively editing. Analysis and ablations are conducted to reveal the query token selection pattern.
[…]
@@ Appendix A Limitations
Due to the limited computational resources, we only extend the model size to 7B and leave the larger model size to future work. We are aware that the original gisting work [gisting] conducts experiments on three model architectures, i.e., encoder-decoder, encoder-only, and decoder-only. In this work, we determine to focus on the decoder-only autoregressive architecture as it is the structure used by most of the popular models nowadays [qwen2, gpt-4, deepseek-v3].
@@
In this work, we maintain the compression rate to roughly around 12:1 (one edit, which contains around 12 tokens for all testing datasets except DUNE [dune], corresponds to one gist token), as one edit represents a fine-grained piece of information. However, we believe it is necessary to investigate the impact of lowering the compression rate, since it potentially helps extend the length of a single edit [silver-bullet].
@@ Appendix B Training details
InComeS is trained on around 1.5 billion tokens, which mainly come from summarization and QA datasets. Specifically, for summerization datasets, we select $4.5e^{6}$ instances from S2ORC [s2orc], $1.15e^{6}$ instances from AG News Corpus  [^8] https://huggingface.co/datasets/sentence-transformers/agnews; and for QA datasets, we use squad [squad], a modified version  [^9] https://huggingface.co/datasets/LLukas22/nq-simplified of the natural question dataset [natural-q], OpenBookQA [openbookqa], QASC [qasc], MedMCQA [medmcqa], and NetEval [neteval]. We also include the training split of ZsRE [zsre], COUNTERFACT [memit], $Wiki_{counterfact}$ from EasyEdit framework [EasyEdit].
We use a cosine linear-warmup scheduler for both models, with a maximum learning rate $1e^{-5}$ and a minimum learning rate $1e^{-6}$ for Llama-3.2-1B and a maximum $5e^{-6}$ and a minimum $1e^{-6}$ for Qwen2.5-7B. To improve the model’s robustness and sample them at a predefined rate during training, the batch size is dynamically sampled from a predefined set rather than a fixed number. Specifically, the predefined set for batch size is 8, 16, 32, 64, and 128, and their corresponding sample rates are 0.05, 0.05, 0.05, 0.15, and 0.7. We adopt DeepSpeed [zero, zerO-infinity, zero-offload] and Liger Kernel [liger-kernel] with 8 Nvidia GPUs for distributed training. Overall, the training takes around 11 hours for Llama-3.2-1B and 35 hours for Qwen2.5-7B.
@@ Appendix C Experiment details
Method Model $\text{WikiData}_{counterfact}$ ZsRE-extended Edit Success Portability Locality Edit Success Portability Locality Base Llama-3.2-1B 21.28 / 21.28 19.73 / 19.73 - 30.06 / 30.06 40.17 / 40.17 - FT-M 97.02 / 94.58 53.43 / 47.51 49.01 / 20.54 99.81 / 95.94 62.80 / 54.84 75.59 / 59.28 LoRA [LoRA] 98.91 / 82.61 52.87 / 43.84 23.31 / 14.94 99.86 / 93.18 57.43 / 44.85 34.94 / 37.41 ROME [rome] 94.33 / 44-33 40.44 / 44-33 26.13 / 44-33 95.41 / 44-33 46.04 / 44-33 40.19 / 44-33 R-ROME [r-rome] 94.31 / 44-33 40.65 / 44-33 25.83 / 44-33 95.29 / 44-33 46.46 / 44-33 39.95 / 44-33 MEMIT [memit] 44-33 / 66.94 44-33 / 23.51 44-33 / 11.12 44-33 / 58.79 44-33 / 25.68 44-33 / 30.10 EMMET [emmet] 44-33 / 27.02 44-33 / 08.86 44-33 / 100.0 44-33 / 15.96 44-33 / 06.84 44-33 / 100.0 MEND [mend] 44-33 / 26.66 44-33 / 21.06 43-33 / 13.54 34-33 / 43.33 34-33 / 30.77 34-33 / 38.75 GRACE [grace] 33.27 / 25.06 14.33 / 10.51 28.71 / 11.43 32.00 / 24.44 12.73 / 10.91 24.12 / 10.12 KN [kn] 20.60 / 34-33 17.16 / 34-33 19.46 / 43-33 16.01 / 43-33 06.70 / 43-33 21.23 / 34-33 IKE [ike] 61.70 / 34-33 45.55 / 34-33 48.80 / 43-33 59.15 / 43-33 57.39 / 43-33 40.21 / 34-33 SERAC [serac] 89.56 / 78.32 60.56 / 46.45 45.67 / 39.36 92.69 / 89.61 66.59 / 63.60 48.96 / 41.32 ICL 93.31 / 82.95 65.81 / 49.75 62.23 / 41.49 68.86 / 60.84 62.19 / 55.58 54.09 / 39.85 InComeS 91.16 / 76.81 65.15 / 45.66 37.35 / 24.64 97.22 / 87.09 70.70 / 52.23 43.91 / 38.97 Base Qwen2.5-7B 22.35 / 22.35 21.46 / 21.46 - 36.21 / 36.21 43.86 / 43.86 - FT-M 98.93 / 90.18 49.39 / 43.13 15.73 / 12.93 99.51 / 92.60 50.04 / 46.41 18.00 / 27.27 LoRA [LoRA] 77.31 / 72.22 37.04 / 31.65 0.380 / 02.24 86.88 / 77.78 28.61 / 24.13 1.290 / 0.330 ROME [rome] 92.69 / 34-33 40.25 / 34-33 38.76 / 34-33 97.86 / 34-33 50.43 / 34-33 51.38 / 34-33 R-ROME [r-rome] 92.59 / 34-33 40.15 / 34-33 38.70 / 34-33 97.89 / 34-33 50.47 / 34-33 51.31 / 34-33 MEMIT [memit] 34-33 / 91.16 34-33 / 39.85 34-33 / 25.85 34-33 / 93.28 34-33 / 49.97 34-33 / 51.85 EMMET [emmet] 34-33 / 88.05 34-33 / 39.37 34-33 / 100.0 34-33 / 93.64 34-33 / 48.44 34-33 / 100.0 MEND [mend] 34-33 / 35.13 34-33 / 15.29 34-33 / 07.29 34-33 / 50.91 34-33 / 38.83 34-33 / 47.12 GRACE [grace] 31.34 / 33.77 25.60 / 18.55 19.19 / 07.29 33.27 / 26.79 14.35 / 11.25 12.31 / 23.45 KN [kn] 36.33 / 34-33 29.60 / 34-33 32.79 / 34-33 14.49 / 34-33 8.49 / 34-33 33.21 / 34-33 IKE [ike] 96.40 / 34-33 75.33 / 34-33 46.98 / 34-33 99.75 / 34-33 83.17 / 34-33 43.19 / 34-33 SERAC [serac] 91.79 / 80.68 51.12 / 41.26 37.84 / 35.10 91.12 / 82.56 62.41 / 52.63 40.56 / 42.46 ICL 90.24 / 85.28 66.99 / 51.66 52.63 / 40.97 71.75 / 71.57 66.10 / 64.57 51.43 / 47.12 InComeS 90.96 / 71.44 66.69 / 47.93 52.04 / 42.02 97.95 / 91.29 75.63 / 61.22 57.16 / 59.21
Table 6: More results for $WikiData_{counterfact}$ [ripple-effects, knowedit] and ZsRE-extended [Yao, knowedit]. The data format of each cell is in "single-edit result / 100-edits result".
@@ C.1 Datasets
The dataset MQuAKE [mquake] (Multi-hop Question Answering for Knowledge Editing) is constructed based on Wikidata and contains question answering instances that require 2-hop, 3-hop, and 4-hop reasoning. In the experiment, we use the latest version of the dataset 101010”MQuAKE-CF-3k-v2.json” in https://github.com/princeton-nlp/MQuAKE, which fixes the knowledge conflict problem for the old version multi-edit subset, and report the accuracy for each query.
@@
DUNE [dune] is a benchmark designed for edits in natural language form. It evaluates the model’s capability of conducting natural language edits through four aspects: scientific reasoning, arithmetic reasoning, new information, and debiasing. As illustrated in Table 2 of [dune], the arithmetic reasoning edits do not follow natural language form as other subsets do and cannot represent a complete piece of instruction, therefore, we do not include it in our experiment.
@@
The $\text{WikiData}_{counterfact}$ [ripple-effects, knowedit] collects triplets from top-viewed pages from Wikipedia and contains portability (ripple-effect [ripple-effects]) instances to test whether the output to the input relevant to the edits is changed as well. Specifically, the portability evaluates the post-edited model from three aspects, including logical generalization, subject aliasing, and reasoning.
@@
The extended version of ZsRE [knowedit, Yao] is constructed based on the original ZsRE [zsre], which is a dataset that focuses on the QA task. The extended version introduces a portability test [Yao], including inverse relation, one-hop reasoning, and subject aliasing.
@@
COUNTERFACT [memit] is a dataset that concentrates on counterfactual information, which typically receives a lower prediction score than accurate facts. It constructs out-of-scope data by substituting the subject entity with a comparable description that has the same predicate.

### v2

@@ 4.3 Ablation study & Analysis
The motivation of including the zero-gist mechanism is to ensure that context-independent tokens can bypass the influence of the edit contexts. To assess the impact of zero-gist, we train a model without this mechanism and evaluate it on DUNE [1] and MQuAKE [49] (see the “w/o zero-gist” line in Table 4). The results show a notable performance drop, particularly in multi-hop scenarios, suggesting that the cross-attention calculations may sometimes interfere with ordinary generation and our zero-gist strategy can mitigate this issue by allowing tokens to “attend to nothing”.
@@
We present the reason for our decision to use the KV cache from the second half of the model layers. To investigate this, we train a model using the KV cache from all layers and evaluate it on 1000 instances from ZsRE [20]. We record the probabilities allocated to the zero-gist in the cross-attention modules, as shown in Fig. 4b. The result shows that the zero-gist probabilities in layers 7-15 are generally lower than those in layers 1-6, and there is a notable drop in the zero-gist probability at layer 7. This suggests that, even when trained to use the full-model KV cache, the model mainly relies on information from deeper layers, since higher probabilities of the zero-gists indicate lower utilization of the actual edit contexts. A possible explanation is that more information is accumulated in the deeper layers, which aids both compression and selection processes. To verify our analysis, we also test the full layer trained model (see the “w/ full model” line in Table 4) under a batch editing setting, and observe a general decrease across all metrics. Additionally, restricting the KV cache to only the second half of the model could provide efficiency benefits with lower memory and computation costs.
@@
Applying a small temperature to the gist cross-attention sharpens the probability distribution over the gist KV caches, which facilitates the model’s ability to retrieve the correct information. We determine the appropriate temperature based on entropy, which has been shown to be an important factor in attention mechanisms [47]. Specifically, we aim to keep the cross-attention entropy close to its optimal value, which occurs when it only needs to attend to one edit. To achieve this, we select 1000 instances from ZsRE [20] and calculate the entropy of the edit batch size 1. We then calculate the entropy for larger edit batch sizes (10, 20, 40, 60, 80, 100, 200, 300) and find the temperatures that align their entropy with the optimal case via gradient descent  [^6] Note that all entropy is calculated in a way that the query and answer part are just a copy of the context. We report the calculated temperature in Figure 4a. As expected, the temperature decreases as the edit batch size increases, but interestingly, it gradually converges to a specific value. Specifically, layers 9, 10, and 14 finally converge to around 0.5. To encourage more decisive selection, we slightly lower this value and set the temperature to $T=0.45$.
@@
Side effect analysis
The side effect analysis is provided in Appendix D.2.
@@ 4.4 Efficiency Analysis
Finally, we present the efficiency analysis for our method. By default, the individual edit length is around 10 to 11. We first compare the efficiency of our method with the efficiency of other knowledge editing methods. Table 5 reports the time required to perform 100 edits for each method. Our method has significantly better efficiency than the other presented editing methods. Additionally, compared to ICL, our approach only needs to maintain the KV cache of the gist representations from the deeper half layers, resulting in substantially lower memory cost. To verify our method’s superiority in efficiency on long context, we further conduct experiments on scaled context length (Table 6). The result demonstrates the efficiency advantage of our method in both the encoding and decoding stages. More detailed analysis can be found in Appendix D.3.
Model InComeS-Compression ICL-Prefilling FT-M LoRA [16] MEMIT [25] EMMET [11] Llama-3.2-1B 0.0326 0.8934 3.4124 33.5412 112.2238 158.6524 Qwen2.5-7B 0.1071 0.9082 28.6132 122.1876 423.4823 512.4235
Table 5: Measured time (seconds) for 100 edits.
Number of edits Model Encoding Decoding InComeS-Compression ICL-Prefilling InComeS-Selection ICL-Generation (with prefilled cache) 1k Llama-3.2-1B 0.2108 1.0413 0.0274 0.1555 2k 0.4051 1.2165 0.0297 0.3545
Table 6: Scaled efficiency comparison (seconds) between InComeS and ICL.
@@ 5 Related Work
The area of knowledge editing (or model editing) has experienced a thriving development in recent years. Researchers have explored various directions in this area. One typical direction is to adopt external memory for the edits. The memory formats applied by different researchers are diverse. Methods like SERAC [29], IKE [48], MeLLo [50] adopt explicit non-parametric memory, which stores specific edit instances, and a retriever that is responsible for recalling relevant edits from the memory. For example, IKE uses KNN, and SERAC applies a trained classifier. Another line of work applies implicit parametric memory. CaliNET [7], T-Patcher [17] embeds the knowledge into a fixed number of neurons and adds them to the model. GRACE [12] adopts a discrete key-value codebook with the value optimized for the desired knowledge. MELO [44] applies dynamic LoRA blocks and indexes them via an internal vector database. KE [4], MEND [28] train a separate meta-model for editing. Another popular direction is to merge knowledge into the model directly. Methods like KN [3], ROME [24], R-ROME [10], MEMIT [25], PMET [22], CoachHooK [21], and AlphaEdit [8] perform editing by tweaking the located FFN part of the model directly. However, some studies reveal that these methods could potentially bring about side effects in the original model [9, 33], leaving the real effectiveness of these methods to be further investigated.
@@ 6 Conclusion
In this paper, we propose InComeS, a scalable model editing method that integrates compression and selection mechanisms directly into the LLMs. InComeS adopts a context compression technique to condense the editing context to KV representations on top of the introduced gist tokens and takes advantage of the compressed KVs to efficiently retrieve the relevant editing context information. Experiments on four different and complex editing settings demonstrate the superiority of our method for comprehensively editing. Further Analysis and ablations validate each component of InComeS and demonstrate the advantage of our method in efficiency.
@@ Appendix A Limitations
Due to the limited computational resources, we only extend the model size to 7B and leave the larger model size to future work. We are aware that the original gisting work [30] conducts experiments on three model architectures, i.e., encoder-decoder, encoder-only, and decoder-only. In this work, we determine to focus on the decoder-only autoregressive architecture as it is the structure used by most of the popular models nowadays [42, 31, 5].
@@
In this work, we maintain the compression rate to roughly around 12:1 (one edit, which contains around 12 tokens for all testing datasets except DUNE [1], corresponds to one gist token), as one edit represents a fine-grained piece of information. However, we believe it is necessary to investigate the impact of lowering the compression rate, since it potentially helps extend the length of a single edit [6].
@@ Appendix B Training details
InComeS is trained on around 1.5 billion tokens, which mainly come from summarization and QA datasets. Specifically, for summerization datasets, we select $4.5e^{6}$ instances from S2ORC [23], $1.15e^{6}$ instances from AG News Corpus  [^7] https://huggingface.co/datasets/sentence-transformers/agnews; and for QA datasets, we use squad [36], a modified version  [^8] https://huggingface.co/datasets/LLukas22/nq-simplified of the natural question dataset [19], OpenBookQA [27], QASC [18], MedMCQA [32], and NetEval [26]. We also include the training split of ZsRE [20], COUNTERFACT [25], $Wiki_{counterfact}$ from EasyEdit framework [39].
We use a cosine linear-warmup scheduler for both models, with a maximum learning rate $1e^{-5}$ and a minimum learning rate $1e^{-6}$ for Llama-3.2-1B and a maximum $5e^{-6}$ and a minimum $1e^{-6}$ for Qwen2.5-7B. To improve the model’s robustness and sample them at a predefined rate during training, the batch size is dynamically sampled from a predefined set rather than a fixed number. Specifically, the predefined set for batch size is 8, 16, 32, 64, and 128, and their corresponding sample rates are 0.05, 0.05, 0.05, 0.15, and 0.7. We adopt DeepSpeed [35, 34, 37] and Liger Kernel [15] with 8 Nvidia GPUs for distributed training. Overall, the training takes around 11 hours for Llama-3.2-1B and 35 hours for Qwen2.5-7B.
@@ Appendix C Experiment details
Method Model $\text{WikiData}_{counterfact}$ ZsRE-extended Edit Success Portability Locality Edit Success Portability Locality Base Llama-3.2-1B 21.28 / 21.28 19.73 / 19.73 - 30.06 / 30.06 40.17 / 40.17 - FT-M 97.02 / 94.58 53.43 / 47.51 49.01 / 20.54 99.81 / 95.94 62.80 / 54.84 75.59 / 59.28 LoRA [16] 98.91 / 82.61 52.87 / 43.84 23.31 / 14.94 99.86 / 93.18 57.43 / 44.85 34.94 / 37.41 ROME [24] 94.33 / 44-33 40.44 / 44-33 26.13 / 44-33 95.41 / 44-33 46.04 / 44-33 40.19 / 44-33 R-ROME [10] 94.31 / 44-33 40.65 / 44-33 25.83 / 44-33 95.29 / 44-33 46.46 / 44-33 39.95 / 44-33 MEMIT [25] 44-33 / 66.94 44-33 / 23.51 44-33 / 11.12 44-33 / 58.79 44-33 / 25.68 44-33 / 30.10 EMMET [11] 44-33 / 27.02 44-33 / 08.86 44-33 / 100.0 44-33 / 15.96 44-33 / 06.84 44-33 / 100.0 MEND [28] 44-33 / 26.66 44-33 / 21.06 43-33 / 13.54 34-33 / 43.33 34-33 / 30.77 34-33 / 38.75 GRACE [12] 33.27 / 25.06 14.33 / 10.51 28.71 / 11.43 32.00 / 24.44 12.73 / 10.91 24.12 / 10.12 KN [3] 20.60 / 34-33 17.16 / 34-33 19.46 / 43-33 16.01 / 43-33 06.70 / 43-33 21.23 / 34-33 IKE [48] 61.70 / 34-33 45.55 / 34-33 48.80 / 43-33 59.15 / 43-33 57.39 / 43-33 40.21 / 34-33 SERAC [29] 89.56 / 78.32 60.56 / 46.45 45.67 / 39.36 92.69 / 89.61 66.59 / 63.60 48.96 / 41.32 ICL 93.31 / 82.95 65.81 / 49.75 52.70 / 45.59 68.86 / 60.84 62.19 / 55.58 62.78 / 59.43 InComeS 91.16 / 76.81 65.15 / 45.66 55.22 / 53.97 97.22 / 87.09 70.70 / 52.23 61.33 / 59.66 Base Qwen2.5-7B 22.35 / 22.35 21.46 / 21.46 - 36.21 / 36.21 43.86 / 43.86 - FT-M 98.93 / 90.18 49.39 / 43.13 15.73 / 12.93 99.51 / 92.60 50.04 / 46.41 18.00 / 27.27 LoRA [16] 77.31 / 72.22 37.04 / 31.65 0.380 / 02.24 86.88 / 77.78 28.61 / 24.13 1.290 / 0.330 ROME [24] 92.69 / 34-33 40.25 / 34-33 38.76 / 34-33 97.86 / 34-33 50.43 / 34-33 51.38 / 34-33 R-ROME [10] 92.59 / 34-33 40.15 / 34-33 38.70 / 34-33 97.89 / 34-33 50.47 / 34-33 51.31 / 34-33 MEMIT [25] 34-33 / 91.16 34-33 / 39.85 34-33 / 25.85 34-33 / 93.28 34-33 / 49.97 34-33 / 51.85 EMMET [11] 34-33 / 88.05 34-33 / 39.37 34-33 / 100.0 34-33 / 93.64 34-33 / 48.44 34-33 / 100.0 MEND [28] 34-33 / 35.13 34-33 / 15.29 34-33 / 07.29 34-33 / 50.91 34-33 / 38.83 34-33 / 47.12 GRACE [12] 31.34 / 33.77 25.60 / 18.55 19.19 / 07.29 33.27 / 26.79 14.35 / 11.25 12.31 / 23.45 KN [3] 36.33 / 34-33 29.60 / 34-33 32.79 / 34-33 14.49 / 34-33 8.49 / 34-33 33.21 / 34-33 IKE [48] 96.40 / 34-33 75.33 / 34-33 46.98 / 34-33 99.75 / 34-33 83.17 / 34-33 43.19 / 34-33 SERAC [29] 91.79 / 80.68 51.12 / 41.26 37.84 / 35.10 91.12 / 82.56 62.41 / 52.63 40.56 / 42.46 ICL 90.24 / 85.28 66.99 / 51.66 52.63 / 40.97 71.75 / 71.57 66.10 / 64.57 51.43 / 47.12 InComeS 90.96 / 71.44 66.69 / 47.93 52.04 / 42.02 97.95 / 91.29 75.63 / 61.22 57.16 / 59.21
Table 7: More results for $WikiData_{counterfact}$ [2, 46] and ZsRE-extended [43, 46]. The data format of each cell is in "single-edit result / 100-edits result".
@@ C.1 Datasets
The dataset MQuAKE [49] (Multi-hop Question Answering for Knowledge Editing) is constructed based on Wikidata and contains question answering instances that require 2-hop, 3-hop, and 4-hop reasoning. In the experiment, we use the latest version of the dataset 999”MQuAKE-CF-3k-v2.json” in https://github.com/princeton-nlp/MQuAKE, which fixes the knowledge conflict problem for the old version multi-edit subset, and report the accuracy for each query.
@@
DUNE [1] is a benchmark designed for edits in natural language form. It evaluates the model’s capability of conducting natural language edits through four aspects: scientific reasoning, arithmetic reasoning, new information, and debiasing. As illustrated in Table 2 of [1], the arithmetic reasoning edits do not follow natural language form as other subsets do and cannot represent a complete piece of instruction, therefore, we do not include it in our experiment.
@@
The $\text{WikiData}_{counterfact}$ [2, 46] collects triplets from top-viewed pages from Wikipedia and contains portability (ripple-effect [2]) instances to test whether the output to the input relevant to the edits is changed as well. Specifically, the portability evaluates the post-edited model from three aspects, including logical generalization, subject aliasing, and reasoning.
@@
The extended version of ZsRE [46, 43] is constructed based on the original ZsRE [20], which is a dataset that focuses on the QA task. The extended version introduces a portability test [43], including inverse relation, one-hop reasoning, and subject aliasing.
@@
COUNTERFACT [25] is a dataset that concentrates on counterfactual information, which typically receives a lower prediction score than accurate facts. It constructs out-of-scope data by substituting the subject entity with a comparable description that has the same predicate.